OpenStudy (anonymous):

Find the points where the graph of y+7=(x-3)^2 crosses the x -axis. The x -coordinate of the right most point is

OpenStudy (pfenn1):

The points where the line crosses the x-axis is found by setting y=0. So\[0+7=(x-3)^2\]Rearrange the formula to solve for x.\[\sqrt{7}=|x-3|\]

OpenStudy (pfenn1):

The right most point would be the highest (most positive) of the two answers.